# AttnLRP: Attention-Aware Layer-wise Relevance Propagation for Transformers

### v1

@@ Abstract
Large Language Models are prone to biased predictions and hallucinations, underlining the paramount importance of understanding their model-internal reasoning process. However, achieving faithful attributions for the entirety of a black-box transformer model and maintaining computational efficiency is an unsolved challenge. By extending the Layer-wise Relevance Propagation attribution method to handle attention layers, we address these challenges effectively. While partial solutions exist, our method is the first to faithfully and holistically attribute not only input but also latent representations of transformer models with the computational efficiency similar to a singular backward pass. Through extensive evaluations against existing methods on Llama 2, Flan-T5 and the Vision Transformer architecture, we demonstrate that our proposed approach surpasses alternative methods in terms of faithfulness and enables the understanding of latent representations, opening up the door for concept-based explanations. We provide an open-source implementation on GitHub [^1] https://github.com/rachtibat/LRP-for-Transformers.
@@ 1 Introduction
To overcome these limitations, it is crucial to understand the latent reasoning process of transformer models. Researchers started using the attention mechanism of transformers as a means to understand how input tokens interact with each other. Attention maps contain rich information about the data distribution Clark et al., (2019); Caron et al., (2021), even allowing for image data segmentation. However, attention, by itself, is inadequate for comprehending the full spectrum of model behavior Wiegreffe and Pinter, (2019). Similar to latent activations, attention is not class-specific and solely provides an explanation for the softmax output (in attention layers) while disregarding other model components. Recent works Geva et al., (2021); Dai et al., (2022), e.g., have discovered that factual knowledge in Large Language Models (LLMs) is stored in Feed-Forward Network (FFN) neurons, separate from attention layers.
@@
Figure 1: By optimizing LRP for transformer-based architectures, our LRP variant outperforms other state-of-the-art methods in terms of explanation faithfulness and computational efficiency. We further are able to explain latent neurons inside and outside the attention module, allowing us to interact with the model. A more detailed discussion on the differences between AttnLRP and other LRP variants can be found in Appendix A.2.2. Heatmaps for other methods are illustrated in Appendix Figure 6. Legend: highly ($+$), semi- ($\circ$), not suited (--). Credit: Nataba/iStock.
@@
Further, attention-based attribution methods such as rollout Abnar and Zuidema, (2020); Chefer et al., 2021a result in checkerboard artifacts, as visible in Figure 1 for a Vision Transformer (ViT). Researchers thus have turned to model-agnostic approaches that aim to provide a holistic explanation of the model’s behavior Miglani et al., (2023), including, e.g., perturbation and gradient-based methods. Methods based on feature perturbation require excessive amounts of compute time (and energy), making their application uneconomically, especially for large architectures. Gradient-based methods, on the other hand, are highly efficient but suffer from noisy gradients and low faithfulness, as evaluated in Section 4.1. Consequently, for understanding latent neurons, perturbation methods remain often the preferred choice Meng et al., (2022); Bills et al., (2023).
@@
Another option is to take advantage of the versatility of rule-based backpropagation methods, such as Layer-wise Relevance Propagation (LRP). These methods allow for the customization of propagation rules to accommodate novel operations, allowing for more faithful explanations and requiring only a singular backward pass. Attempts to apply LRP to non-linear attention, however, have faced challenges such as numerical instability, lack of faithfulness, or completely circumventing the non-linear softmax problem, see Appendix A.2.2.
@@
In this work, we introduce AttnLRP, an extension of LRP within the Deep Taylor Decomposition framework Montavon et al., (2017), with the particular requirements necessary for attributing non-linear attention accurately. AttnLRP allows explaining transformer-based models with high faithfulness and efficiency, while also allowing attribution of latent neurons and providing insights into their role in the generation process (see Figure 2).
@@
    We illustrate how to gain insights into a LLM generation process by identifying relevant neurons and explaining their encodings.
@@ 2.1 Perturbation & Local Surrogates
In perturbation analysis, such as occlusion-based attribution Zeiler and Fergus, (2014) or SHAP Lundberg and Lee, (2017), the input features are repeatedly perturbed while the effect on the model output is measured Fong and Vedaldi, (2017). AtMan Deb et al., (2023) is specifically adapted to the transformer architecture, where tokens are not suppressed in the input space, but rather in the latent attention weights. Interpretable local surrogates, on the other hand, replace complex black-box models with simpler linear models that locally approximate the model function being explained. Since the surrogate has low complexity, interpretability is facilitated. Prominent methods include LIME Ribeiro et al., (2016) and LORE Guidotti et al., (2018).
@@ 2.2 Attention-based
These methods take advantage of the attention mechanism in transformer models. Although attention maps capture parts of the data distribution, they lack class specificity and do not provide a meaningful interpretation of the final prediction Wiegreffe and Pinter, (2019). Attention Rollout Abnar and Zuidema, (2020) attempts to address the issue by sequentially connecting attention maps of all layers. However, the resulting attributions are still not specific to individual outputs and exhibit substantial noise. Hence, Gildenblat, (2023) has found that reducing noise in attention rollout can be achieved by filtering out excessively strong outlier activations. To enable class-specificity, the work of Chefer et al., 2021b proposed a novel rollout procedure wherein the attention’s activation is mean-weighted using a combination of the gradient and LRP-inspired relevances. It is important to note that this approach yields an approximation of the mean squared relevance value, which diverges from the originally defined notion of ”relevance“ or ”importance“ of additive explanatory models such as SHAP Lundberg and Lee, (2017) or LRP Bach et al., (2015). Subsequent empirical observations by Chefer et al., 2021a revealed that an omission of LRP-inspired relevances and a sole reliance on a positive mean-weighting of the attention’s activation with the gradient improved the faithfulness inside cross-attention layers. Though, this approach can only attribute positively and does not consider counteracting evidence.
Attention-rollout based approaches, while offering advantages in terms of computational efficiency and conceptual simplicity, have significant drawbacks. Primarily, they suffer from a limited resolution in the input attribution maps, resulting in undesirable checkerboard artifacts cf. Figure 1. Moreover, they are unable to attribute hidden latent features beyond the softmax output. Consequently, these approaches only provide explanations for a fraction of the model, thereby compromising the overall fidelity and limiting the feasibility of explanations within the hidden space.
@@ 2.3 Backpropagation-based
Modified backpropagation methods, such as LRP Bach et al., (2015), decompose individual layers instead of linearizing the entire model. They modify the gradient to produce more reliable attributions Arras et al., (2022). The work Ding et al., (2017) was the first to apply standard LRP on non-linear attention layers, while Voita et al., (2021) proposed an improved variant building upon Deep Taylor decomposition. Nonetheless, both variants can lead to numerical instabilities in the softmax and do not fulfill the conservation property (3) in matrix multiplication (Lemma 3 of Chefer et al., 2021b ). Ali et al., (2022) considerably improved attributions by regarding the non-linear softmax operation as constant attributing relevance solely through the value path and disregarding the softmax operation entirely. In Appendix A.2.2, we compare the different LRP-variants.
@@ 3.1 Layer-wise Relevance Propagation
This provides us with ”importance values” for the input variables, which reveal their direct contribution to the final prediction.
Unlike other methods, LRP treats a neural network as a layered directed acyclic graph, where each neuron $j$ in layer $l$ is modeled as a function node $f^{l}_{j}$ that is individually decomposed according to equation (1). Beginning at the model output $L$, the initial relevance value $R^{L}_{j}\propto f^{L}_{j}$ is successively distributed to its prior network neurons one layer at a time. Hence, LRP follows the flow of activations computed during the forward pass through the model in the opposite direction, from output $f^{L}$ back to input layer $f^{1}$.
@@ 3.1.1 Decomposition through Linearization
where $\mathcal{O}$ is the approximation error in Big-$\mathcal{O}$ notation and the Jacobian J is evaluated at reference point $\tilde{\textbf{x}}$, that is in the following omitted for brevity222if $=\textbf{x}$ x, this is equivalent to Gradient $\times$ Input. The bias term represents the constant portion of the function and the approximation error that cannot be directly attributed to the input variables.
We substitute the neural network function with its first-order expansion and assert its proportionality to a relevance value $R_{j}$ through multiplication with a constant factor $c\in\mathbb{R}$ with $f_{j}(\textbf{x})\neq 0$.
$R_{j}=f_{j}(\textbf{x})\ c=\left(\sum_{i}\textbf{J}_{ji}\ x_{i}+\tilde{b}_{j}\right)\frac{R_{j}}{f_{j}(\textbf{x})}$
Generally, we focus on attributing the input variables and hence ignore the contribution of the bias term Bach et al., (2015). However, it is important to note that the bias absorbs part of the relevance and that the conservation property (3) still holds if the bias is regarded as an additional neuron without connection to the input. Alternatively, the bias term can be distributed equally among the input variables, as explained in Appendix A.2.1.
Comparing with Equation (1), we identify $R_{i\leftarrow j}$ as the summands and apply Equation (2) to the input variables. In addition, we insert stabilizing factor $\varepsilon$ (usually set to $10^{-6}$) with the sign of $f_{j}(\textbf{x})$ to allow for the case $f_{j}(\textbf{x})=0$:
@@
In the following, $\text{sign}(f_{j}(\textbf{x}))$ is omitted for brevity.
@@
$\Rightarrow\textbf{R}^{l-1}=\textbf{x}\odot\textbf{J}^{T}\cdot\textbf{R}^{l}\oslash(\textbf{f}(\textbf{x})+\varepsilon)$
@@ 3.2.1 The $\varepsilon$- and $\gamma$-LRP rule
$R_{i}=\sum_{j}\textbf{W}_{ji}x_{i}\frac{R_{j}}{z_{j}(\textbf{x})+\varepsilon}$ (8)
The bias $b_{j}$ (Equation (6)) absorbs a portion of the relevance. The proof is omitted for brevity. We employ the $\varepsilon$-LRP rule on all linear layers, unless specified otherwise.
In models with many layers, the gradient of a linear layer (equal to the weight matrix W) can cause noisy attributions due to the gradient shattering effect Balduzzi et al., (2017); Dombrowski et al., (2022). To mitigate this noise, it is best practice to use the $\gamma$-LRP rule Montavon et al., (2019), an extension to improve the signal-to-noise ratio. We have observed that this effect is significantly pronounced in ViTs while LLMs lack visible noise. Therefore, we only apply the $\gamma$-LRP rule to linear layers in ViTs. For more details, please refer to Appendix A.2.3.
@@ 3.3 Attributing Non-linear Attention
$\textbf{A}=\text{softmax}\left(\frac{\textbf{Q}\cdot\textbf{K}^{T}}{\sqrt{d_{k}}}\right)$ (10)
@@
where ($\cdot$) denotes matrix multiplication, $\textbf{K}\in\mathbb{R}^{b\times k\times d_{k}}$ is the key matrix, $\textbf{Q}\in\mathbb{R}^{b\times q\times d_{k}}$ is the queries matrix, and $\textbf{V}\in\mathbb{R}^{b\times k\times d_{v}}$ the values matrix, and $\textbf{O}\in\mathbb{R}^{b\times k\times d_{v}}$ is the final output of the attention mechanism. $b$ is the batch dimension including the number of heads, and $d_{k},d_{v}$ indicate the embedding dimensions, and $s_{q},s_{k}$ are the number of query and key/value tokens.
@@ 3.3.1 Handling the Softmax Non-Linearity
In Section 3.1.1, we present a generalized approach to linearization that incorporates bias terms, allowing for the absorption of a portion of the relevance. However, Ali et al., (2022) advocates for a strict adherence to the conservation property (3) and argues that a linear decomposition of a non-linear function should typically exclude a bias term. While we see the virtue of this approach for operations such as RMSNorm Zhang and Sennrich, (2019) or matrix multiplication, where $f(0)=0$, we contend that a linearization of the softmax function should inherently incorporate a bias term. This is due to the fact that even when the input is zero, the softmax function yields a value of $\frac{1}{N}$ (where $N$ represents the dimension of the inputs) which is analogous to a bias term.
@@
$R^{l}_{i}=x_{i}(R^{l+1}_{i}-s_{i}\sum_{j}R^{l+1}_{j})$ (13)
where $s_{j}$ denotes the $j$-th output of the softmax function. The hidden bias term, which represents the approximation error, consequently absorbs a portion of the relevance.
The proof can be found in Appendix A.3.1.
Note, that the works Voita et al., (2021); Chefer et al., 2021b ; Ali et al., (2022) propose to handle the bias term differently to strictly enforce the conservation property (3). Voita et al., (2021) linearizes at x but distributes the bias term equally on all input variables, while Chefer et al., 2021b applies the element-wise identity rule (9) and hence omits the bias term completely. Both variants can lead to severe numerical instabilities as discussed in Appendix A.2.1 and seen empirically in our preliminary experiments. Finally, Ali et al., (2022) regards the attention matrix A in equation (11) as constant, attributing relevance solely through the value path and disregarding the softmax operation entirely. A more detailed discussion on the differences of the LRP variants can be found in Appendix A.2.2.

### v2

@@ Abstract
Large Language Models are prone to biased predictions and hallucinations, underlining the paramount importance of understanding their model-internal reasoning process. However, achieving faithful attributions for the entirety of a black-box transformer model and maintaining computational efficiency is an unsolved challenge. By extending the Layer-wise Relevance Propagation attribution method to handle attention layers, we address these challenges effectively. While partial solutions exist, our method is the first to faithfully and holistically attribute not only input but also latent representations of transformer models with the computational efficiency similar to a single backward pass. Through extensive evaluations against existing methods on LLaMa 2, Mixtral 8x7b, Flan-T5 and vision transformer architectures, we demonstrate that our proposed approach surpasses alternative methods in terms of faithfulness and enables the understanding of latent representations, opening up the door for concept-based explanations. We provide an LRP library at https://github.com/rachtibat/LRP-eXplains-Transformers.
@@ 1 Introduction
To overcome these limitations, it is crucial to understand the latent reasoning process of transformer models. Researchers started using the attention mechanism of transformers as a means to understand how input tokens interact with each other. Attention maps contain rich information about the data distribution Clark et al., (2019); Caron et al., (2021), even allowing for image data segmentation. However, attention, by itself, is inadequate for comprehending the full spectrum of model behavior Wiegreffe and Pinter, (2019). Similar to latent activations, attention is not class-specific and solely provides an explanation for the softmax output (in attention layers) while disregarding other model components. Recent works Geva et al., (2021); Dai et al., (2022) have in fact discovered that factual knowledge in Large Language Models (LLMs) is stored in Feed-Forward Network (FFN) neurons, separate from attention layers.
@@
Figure 1: By optimizing LRP for transformer-based architectures, our LRP variant outperforms other state-of-the-art methods in terms of explanation faithfulness and computational efficiency. We further are able to explain latent neurons inside and outside the attention module, allowing us to interact with the model. A more detailed discussion on the differences between AttnLRP and other LRP variants can be found in Appendix A.2.2. Heatmaps for other methods are illustrated in Appendix Figure B.6. Legend: highly ($+$), semi- ($\circ$), not (--) suited. Credit: Nataba/iStock.
@@
Further, attention-based attribution methods such as rollout Abnar and Zuidema, (2020); Chefer et al., 2021a result in checkerboard artifacts, as visible in Figure 1 for a Vision Transformer (ViT). Researchers thus have turned to model-agnostic approaches that aim to provide a holistic explanation of the model’s behavior Miglani et al., (2023), including, e.g., perturbation and gradient-based methods.
Methods based on feature perturbation require excessive amounts of compute time (and energy), and in order to access latent attributions they require performing perturbations at each layer separately, resulting in further exponential cost increase. This makes their application economically infeasible, especially for large architectures. In contrast, gradient-based methods benefit from the chain-rule in automatic differentiation and can produce latent attributions for all layers in a single backward pass. While prominent gradient-based methods, e.g. Input $\times$ Gradient Simonyan et al., (2014), are highly efficient, they suffer from noisy gradients and low faithfulness, as evaluated in Section 4.1.
@@
Another option is to take advantage of the versatility of rule-based backpropagation methods, such as Layer-wise Relevance Propagation (LRP). These methods allow for the customization of propagation rules to accommodate novel operations, allowing for more faithful explanations and requiring only a single backward pass. As thoroughly discussed in Appendix A.2.2, all previous attempts to apply LRP to transformers reused standard LRP rules Ding et al., (2017); Voita et al., (2021); Chefer et al., 2021b ; Ali et al., (2022). However, transformer architectures include several functions for which standard LRP rules do not adequately apply, such as softmax, bi-linear (matrix) multiplication (e.g. query-key multiplication) and layer normalization. Additionally, the routing networks in Mixture of Experts (MoE) models Fedus et al., (2022) present notable challenges due to their combination of these functions. As a result, other previous attempts result in either numerical instabilities or low faithfulness.
Our method represents a significant breakthrough in handling the attribution problem in transformer architectures by enabling an accurate attribution flow through non-linear model components outperforming other existing methods (including perturbation) by a large margin.
@@
In this work, we introduce AttnLRP, an extension of LRP within the Deep Taylor Decomposition framework Montavon et al., (2017), with the particular requirements necessary for attributing non-linear transformer components accurately. AttnLRP allows explaining transformer-based models with high faithfulness and efficiency, while also allowing attribution of latent neurons and providing insights into their role in the generation process (see Figure 2).
@@
    We illustrate how to gain insights into an LLM generation process by identifying relevant neurons and explaining their encodings.
@@ 2.1 Perturbation & Local Surrogates
In perturbation analysis, such as occlusion-based attribution Zeiler and Fergus, (2014) or SHAP Lundberg and Lee, (2017), the input features are repeatedly perturbed while the effect on the model output is measured Fong and Vedaldi, (2017). AtMan Deb et al., (2023) is specifically adapted to the transformer architecture, where tokens are not suppressed in the input space, but rather in the latent attention weights.
Interpretable local surrogates, on the other hand, replace complex black-box models with simpler linear models that locally approximate the model function being explained. Since the surrogate has low complexity, interpretability is facilitated. Prominent methods include LIME Ribeiro et al., (2016) and LORE Guidotti et al., (2018).
@@ 2.2 Attention-based
These methods take advantage of the attention mechanism in transformer models. Although attention maps capture parts of the data distribution, they lack class specificity and do not provide a meaningful interpretation of the final prediction Wiegreffe and Pinter, (2019). Attention Rollout Abnar and Zuidema, (2020) attempts to address the issue by sequentially connecting attention maps of all layers. However, the resulting attributions are still not specific to individual outputs and exhibit substantial noise. Hence, Gildenblat, (2023) has found that reducing noise in attention rollout can be achieved by filtering out excessively strong outlier activations.
To enable class-specificity, the work of Chefer et al., 2021b proposed a novel rollout procedure wherein the attention’s activation is mean-weighted using a combination of the gradient and LRP-inspired relevances. It is important to note that this approach yields an approximation of the mean squared relevance value, which diverges from the originally defined notion of “relevance” or “importance” of additive explanatory models such as SHAP Lundberg and Lee, (2017) or LRP Bach et al., (2015). Subsequent empirical observations by Chefer et al., 2021a revealed that an omission of LRP-inspired relevances and a sole reliance on a positive mean-weighting of the attention’s activation with the gradient improved the faithfulness inside cross-attention layers. Though, this approach can only attribute positively and does not consider counteracting evidence.
Attention-rollout based approaches, while offering advantages in terms of computational efficiency and conceptual simplicity, have significant drawbacks. Primarily, they suffer from a limited resolution in the input attribution maps, resulting in undesirable checkerboard artifacts cf. Figure 1. Moreover, they are unable to attribute hidden latent features beyond the softmax output. Consequently, these approaches only provide explanations for a fraction of the model, thereby compromising the fidelity and limiting the feasibility of explanations within the hidden space.
@@ 2.3 Backpropagation-based
Modified backpropagation methods, such as LRP Bach et al., (2015), decompose individual layer functions instead of linearizing the entire model. They modify the gradient to produce more reliable attributions Arras et al., (2022). The work Ding et al., (2017) was the first to apply standard LRP on non-linear attention layers, while Voita et al., (2021) proposed an improved variant building upon the Deep Taylor Decomposition framework. Nonetheless, both variants can lead to numerical instabilities in attributing the softmax function and do not fulfill the conservation property (3) in matrix multiplication. Ali et al., (2022) considerably improved attributions by recognizing that standard LRP rules were not suitable for these operations and proposed to exclude softmax and normalization operations from the computational graph by stopping the relevance (gradient) flow through them. However, it does not resolve the fundamental challenge of optimally applying LRP to non-linear operations. In Appendix A.2.2, we provide a comprehensive analysis about different LRP-variants.
@@ 3.1 Layer-wise Relevance Propagation
This provides us with “importance values” for the input variables, which reveal their direct contribution to the final prediction. Unlike other methods, LRP treats a neural network as a layered directed acyclic graph, where each neuron $j$ in layer $l$ is modeled as a function node $f^{l}_{j}$ that is individually decomposed according to Equation (1). Beginning at the model output $L$, the initial relevance value $R^{L}_{j}\propto f^{L}_{j}$ is successively distributed to its prior network neurons one layer at a time. Hence, LRP follows the flow of activations computed during the forward pass through the model in the opposite direction, from output $f^{L}$ back to input layer $f^{1}$.
@@ 3.1.1 Decomposition through Linearization
where $\mathcal{O}$ is the approximation error in Big-$\mathcal{O}$ notation and the Jacobian J is evaluated at reference point $\tilde{\textbf{x}}$, that is in the following omitted for brevity111if $=\textbf{x}$ x, this is equivalent to Gradient $\times$ Input. We have taken the DTD perspective to highlight the bias term, which is important for the upcoming discussion.. The bias term represents the constant portion of the function and the approximation error that cannot be directly attributed to the input variables.
We substitute the layer function with its first-order expansion and assert its proportionality to a relevance value $R_{j}$ following Equation (1) through multiplication with a constant factor $c\in\mathbb{R}$ with $f_{j}(\textbf{x})\neq 0$.
$R_{j}=f_{j}(\textbf{x})\ c=\sum_{i}\underbrace{\textbf{J}_{ji}\ x_{i}\frac{R_{j}}{f_{j}(\textbf{x})}}_{R_{i\leftarrow j}}+\underbrace{\tilde{b}_{j}\frac{R_{j}}{f_{j}(\textbf{x})}}_{R_{b\leftarrow j}}$
Comparing with Equation (1), we identify $R_{i\leftarrow j}$ as the relevance assigned to the input variables and $R_{b\leftarrow j}$ as the relevance assigned to the bias term. Hence, the bias term absorbs a portion of the relevance $R_{j}$ that is not allocated to the input variables. This technically violates the conservation property (3), as only $R_{i\leftarrow j}$ is further distributed to prior layers reducing the amount of relevance per distribution step. However, Bach et al., (2015) treats bias terms as additional hidden neurons (with an activation value of one and a weight that equals the bias value, connected to the output) including them into the conservation property (3). Consequently, we regard this relevance as preserved, rather than lost. Alternatively, to strictly enforce conservation, the absorbed relevance score of the bias term can be distributed equally among the input variables, or the bias term can be excluded completely, as explained in Appendix A.2.1.
To obtain a propagation rule for the input variables, we apply Equation (2) without the bias term. In addition, we insert a stabilizing factor $\varepsilon\ll|f_{j}(\textbf{x})|\in\mathbb{R}^{+}$ with the sign of $f_{j}(\textbf{x})$ to allow for the case $f_{j}(\textbf{x})=0$:
@@
In the following, $\text{sign}(f_{j}(\textbf{x}))\in\{-1,1\}$ is omitted for brevity. Note, that $\varepsilon$ acts as bias term and absorbs a negligible amount of the relevance.
@@
$\Rightarrow\textbf{R}^{l-1}=\textbf{x}\odot\textbf{J}^{\top}\cdot\textbf{R}^{l}\oslash(\textbf{f}(\textbf{x})+\varepsilon)$
@@ 3.2.1 The $\varepsilon$- and $\gamma$-LRP rule
$R_{i}^{l-1}=\sum_{j}\textbf{W}_{ji}x_{i}\frac{R_{j}^{l}}{z_{j}(\textbf{x})+\varepsilon}$ (8)
The bias $b_{j}$ of Equation (6) and $\varepsilon$ absorb a portion of the relevance. The proof is omitted for brevity. We employ the $\varepsilon$-LRP rule on all linear layers, unless specified otherwise.
In models with many layers, the gradient of a layer can cause noisy attributions due to the gradient shattering effect Balduzzi et al., (2017); Dombrowski et al., (2022). To mitigate this noise, it is best practice to use the $\gamma$-LRP rule Montavon et al., (2019), an extension to improve the signal-to-noise ratio. We have observed that this effect is significantly pronounced in ViTs while LLMs lack visible noise. Therefore, we only apply the $\gamma$-LRP rule to linear layers in ViTs. For more details, please refer to Appendix A.2.3.
@@ 3.3 Attributing Non-linear Attention
$\textbf{A}=\text{softmax}\left(\frac{\textbf{Q}\cdot\textbf{K}^{\top}}{\sqrt{d_{k}}}\right)$ (10)
@@
where ($\cdot$) denotes matrix multiplication, $\textbf{K}\in\mathbb{R}^{b\times s_{k}\times d_{k}}$ is the key matrix, $\textbf{Q}\in\mathbb{R}^{b\times s_{q}\times d_{k}}$ is the queries matrix, $\textbf{V}\in\mathbb{R}^{b\times s_{k}\times d_{v}}$ the values matrix, and $\textbf{O}\in\mathbb{R}^{b\times s_{k}\times d_{v}}$ is the final output of the attention mechanism. $b$ is the batch dimension including the number of heads, and $d_{k},d_{v}$ indicate the embedding dimensions, and $s_{q},s_{k}$ are the number of query and key/value tokens.
@@ 3.3.1 Handling the Softmax Non-Linearity
In Section 3.1.1, we present a generalized approach to linearization that incorporates bias terms, allowing for the absorption of a portion of the relevance. However, Ali et al., (2022) advocates for a strict adherence to the conservation property (3) and argues that a linear decomposition of a non-linear function should typically exclude a bias term. While we see the virtue of this approach for operations such as RMSNorm Zhang and Sennrich, (2019) or matrix multiplication, where $f(0)=0$, we contend that a linearization of the softmax function should inherently incorporate a bias term. This is due to the fact that even when the input is zero, the softmax function yields a value of $\frac{1}{N}$ (where $N$ represents the dimension of the inputs) which is analogous to a virtual bias term.
@@
$R^{l-1}_{i}=x_{i}(R^{l}_{i}-s_{i}\sum_{j}R^{l}_{j})$ (13)
where $s_{i}$ denotes the $i$-th output of the softmax function. The hidden bias term, which contains the approximation error, consequently absorbs a portion of the relevance.
The proof can be found in Appendix A.3.1. In Appendix A.2.4, we explore the implications of vanishing gradients and temperature scaling on attributing the softmax function, which is important when attributing softmax outside of the attention mechanism, e.g. at the classification output. Note, that the works Ding et al., (2017); Voita et al., (2021); Chefer et al., 2021b ; Ali et al., (2022) propose to handle the bias term differently to strictly enforce the conservation property (3). Most variants can lead to severe numerical instabilities as discussed in Appendix A.2.1 and seen empirically in our preliminary experiments.